Impact of Increased Sales on Operating Income Using the Degree of Operating Leverage

Chillmax Company had planned to sell 3,500 pairs of shoes at $60 each in the coming year. Unit variable cost is $21 (includes direct materials, direct labor, variable factory overhead, and variable selling expense). Total fixed cost equals $78,000 (includes fixed factory overhead and fixed selling and administrative expense). Operating income at 3,500 units sold is $58,500. The degree of operating leverage is 2.3. Now Chillmax expects to increase sales by 10% next year.

Required:

1. Calculate the percent change in operating income expected.

%

2. Calculate the operating income expected next year using the percent change in operating income calculated in Requirement 1.
Answer #1

Contribution margin = (60-21)*3500 = 136,500

Net operating income = 58,500

Degree of operating leverage = Contribution margin/Net operating income = 136,500/58,500 = 2.33

A) percentage change in operating income

= 10%*2.33

= 23.3% (rounded)

B)

Sales (3500+10%)*60 231000
Variable cost (3500+10%)*21 80,850
Contribution margin 150,150
Fixed cost 78,000
Net operating income 72,150

Note :

I have rounded some numbers. Comment if you face any issues